OpenStudy (anonymous):

Four times the sum of a number and six is the same as the difference of ten and three times the number.find the number

OpenStudy (akashdeepdeb):

Let the 'number' be = x Four times the sum of a number and six = 4 * (x + 6) Difference of 10 and 3 times the number = 3*x - 10 Can you solve the equation now?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

Is it -54?

OpenStudy (akashdeepdeb):

How did you get that?

OpenStudy (anonymous):

4(x+6)=3(x-10) 4x+24=3x-30 Then I subtracted 3x from both sides and got x (4x-3x) then subtracted 24 from both sides so -30-24 so I got x =-54

OpenStudy (akashdeepdeb):

Look at your equations and my equations. \[HINT: ~~3x- 10 ~\neq ~3(x-10)\]

OpenStudy (anonymous):

Ok is it -34

OpenStudy (akashdeepdeb):

Yes. :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

@robtobey can u come help me

OpenStudy (anonymous):

please

OpenStudy (anonymous):

The attachment represents my interpretation of the problem statement.